Is this statement true or false and please explain why:

There is a 9 M aqueous HCl solution in the stock room, but a 4.5 M solution is required for an experiment. Doubling the volume of the 9 M sample with water will produce the 4.5 M solution.
Answer #1

yes it is correct

by dilution formula M1V1 =M2V2

now M1=9M , let V1=1L ,if we double the volume of solution V2=2L

So 9M*1L =M2*2L

So M2=(9M*1L)/2L =4.5M

So doubling the volume of the 9 M sample with water will produce the 4.5 M solution

thank you